OpenStudy (anonymous):

What is 120 degrees in radians?

OpenStudy (anonymous):

2pi/3

OpenStudy (anonymous):

multiply 120 by pi/180, you get 120pi/180 = 2pi/3

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

welcome.

OpenStudy (anonymous):

\[120^0 = 120^0 \times \frac{\pi}{180^0} = \frac{2\pi^c}{3} \] @bridgetclancy

OpenStudy (anonymous):

What is 2 pi/5 radians in degrees?